OpenStudy (anonymous):

A fruit stand has 28 apples, 20 oranges, and 12 pears. What is the ratio of apples to the total pieces of fruit? A. 1 : 4 B. 7 : 15 C. 1 : 2 D. 7 : 8

OpenStudy (amoodarya):

\[\frac{ 28 }{28+20+12 }\]

OpenStudy (amoodarya):

simplify then you will find the answer

OpenStudy (anonymous):

how to simplify 0.4666666666666667

OpenStudy (amoodarya):

\[\frac{ 28 }{ 60}=\frac{ 7*4 }{ 4*15}=\frac{ 7 }{ 15}\]

OpenStudy (amoodarya):

now its correct

OpenStudy (anonymous):

tnx